OpenStudy (anonymous):

Type this equation in standard form. 8y + 4 = -3x @x1MxBATMANx @xoya

OpenStudy (anonymous):

I guess 3x+8y+4=0

hartnn (hartnn):

Ax+By =C

hartnn (hartnn):

so, bring -3x on left and +4 on right...

OpenStudy (anonymous):

So @hartnn I don't think so :/

OpenStudy (anonymous):

how is tht @hartnn

hartnn (hartnn):

the standard form is Ax+By =C and it can be written as Ax+By+C=0 also :) so, 3x+8y+4=0 is also correct :)

OpenStudy (anonymous):

Ohh okay thanks for clearing :)

OpenStudy (anonymous):

3x+8y+4=0 i tried that and it didnt let me

hartnn (hartnn):

then try 3x+8y=-4

OpenStudy (anonymous):

LOL

OpenStudy (anonymous):

Solve this system of equations using the addition method. -8x + 3y = -5 8x - 2y = 6

hartnn (hartnn):

umm...then add those 2 equations...

OpenStudy (anonymous):

i dont get them

hartnn (hartnn):

what u got after adding ?

OpenStudy (anonymous):

idk

hartnn (hartnn):

what is -8x+8x =... ?

OpenStudy (anonymous):

0

hartnn (hartnn):

what is 3y-2y =... ?

OpenStudy (anonymous):

1

hartnn (hartnn):

1 what ? only 1? or 1y ?

OpenStudy (anonymous):

1y

hartnn (hartnn):

yes, whats -5+6 =... ?

OpenStudy (anonymous):

1

hartnn (hartnn):

yes, so adding these -8x + 3y = -5 8x - 2y = 6 gives you 1y = 1 y=1 got this ? can u find x ?

OpenStudy (anonymous):

no i cant find x

OpenStudy (anonymous):

message me @xoya

OpenStudy (anonymous):

What? :/